OpenStudy (anonymous):

Find the y-intercept of the tangent line at the point (-2, f(-2)) on the graph of. f(x)=4x^2+3x+4

OpenStudy (anonymous):

when x=-2 , y= 4(4)-6+4 = 14 so the point is (-2,14). f'(x) = 8x +3 when x=-2, f'(-2) = -13

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y-14= -13(x-(-2) )

OpenStudy (anonymous):

y= -13x -12

OpenStudy (anonymous):

y intercept is (0,-12)

OpenStudy (anonymous):

Thank you!